# Persistence diagrams as morphological signatures of cells: A method to measure and compare cells within a population

### v2

@@ Abstract
Quantifying cellular morphology is a common approach in life sciences to study regulatory mechanisms, cell fate, homeostasis, and pathologies. However, inherent heterogeneities within cell populations can render traditional automated microscopy image analysis methods unreliable or even impossible. This often leads to manual inspection and pseudo-quantitative evaluation of cell images, which is prone to user bias and not readily reproducible. To address these challenges, we developed a computational framework based on topological data analysis (TDA) to investigate and quantify cell morphology heterogeneity with two sample sets exemplifying common cases in cell biology: human adult mesenchymal stem cells, which are primary cells derived from a patient, and HeLa cells, an immortalised cell line. Our approach is based on the characterisation of the shape of the cell by its contour and the position of its nucleus. We build a filtering of the edges defining the contour using a radial distance function initiated from the nucleus. This filtering is then used to construct a persistence diagram that serves as a signature of the cell shape. Two cells can then be compared by computing the Wasserstein distance between their persistence diagrams. Given a cell population, we then compute a distance matrix that includes all pairwise distances between its members. We analyse this distance matrix using hierarchical clustering with different linkage schemes and define a purity score that quantifies consistency between those different schemes, which can then be used to assess homogeneity within the cell population. We illustrate and validate our approach by identifying sub-populations in human mesenchymal stem cell populations, and analysing the heterogeneity of a population of HeLa cells.
@@
The fundamental units of life are cells, and understanding how they grow, divide, die, and change shape is of central importance in many areas of the life sciences. In this paper, we focus on how to define and compare the shape of cells. We identify the shape of a cell with its contour (i.e. location of its membrane) supplemented with the position of its nucleus. We use topological data analysis to define a signature of that shape, generated from its persistent homology, a mathematical construction that reflects the relative position of the nucleus with respect to segments of the contours. We compute the distance between two cells as the Wasserstein distance between their shape signatures. Using this distance, we analyse populations of cells to help identify members with unusual shapes (usually referred to as outliers) as well as sub-populations. We validate our approach to identify sub-populations within human mesenchymal stem cell populations that are known to be heterogeneous, and search for outliers in a population of HeLa cells.
@@ 1 Introduction
A second representation of 2D shapes, which we refer to as shape as planar contour, is based on the curve describing the outer boundary of the shape (see Figure 1C). This is well suited to applications focused on the geometric configuration of a shape, where factors such as the colour or grey level of the interior are either not relevant or not available. Methods for modelling the similarity between two shapes given as planar contours have been based on defining a distance between two curves in the plane. The proposed distances include the Hausdorff and Frechet distances [8]. Other techniques are based on the Poisson equation [9], integral invariants [10], and an elastic shape distance on the energy required to elastically deform one boundary contour to the other [11, 12].
@@
The paper is organised as follows. The next section (Section 2) introduces the concept of persistence homology (section 2.1) applied to analysing the morphology of a cell, the construction of the persistence diagram of a cell contour (section 2.2), and the computation of the Wasserstein distance between two persistence diagrams (section 2.4). The Materials and Methods section (section 3) gives information on the experimental data and implementations of the methods mentioned above. The Results section (section 4) discusses the applications of this new method on two systems, hMSCs (section 4.1) that are expected to include sub-populations and, in contrast, HeLa cells (section 4.2) whose morphologies are expected to vary continuously due to local conditions. We conclude with a discussion (section 5) of how to integrate persistent homology with other more traditional techniques for comparing cell shapes.
@@ 2.3 Illustration on a simple cell contour
Readers familiar with persistent homology and persistence diagrams will notice that this is a non-standard modification. Due to the nature of the contours, performing this essential pairing allows us to more efficiently represent and compare the topological descriptors.
@@ 3.2 HeLa cells
3.3 Fluoresecence Staining of Cells
@@ 3.5 Image Processing and Contour Generation
We used the FilamentSensor2.0 tool [22] to perform image processing and extract the contour of each cell from the actin image. Here, we used the feature “Include Area-Outline” to export the contour from the binarised image of the cells. The centre of the cell is obtained from the centre of mass from the aligned microscopy image of the nucleus, where we thresholded the nucleus in Fiji [23] using the “Otsu” method, before outlining it and determining the $x$- and $y$-coordinates of the centre of mass.
@@ 4 Results and discussion
In this section, we report how our method for comparing the shapes of cells can be used to identify unusual cells (i.e., outliers) and subpopulations within a cell population. We consider both human mesenchymal stem cells (hMSCs) and HeLa cells. HeLa cells are a long-established, immortalised cancer cell line. They are a commonly used cell line for culturing in in vitro conditions, showing relatively consistent growth patterns and morphologies within a population, despite some heterogeneity. hMSCs are primary cells, on the other hand, derived from bone marrow directly from patients. They are not immortalised or further altered and exhibit an inherent natural heterogeneity, both in morphology and function, due to their preparation and selection method.
We analyse a population of hMSCs, $X1$ (Section 4.1) and a population of HeLa cells, $Y1$ (Section 4.2), with the experimental setup and analysis pipeline described in Section 3.
4.1 $X1$: sub-populations of hMSCs
@@
Linkage A (red, $n=86$) B (blue, $n=7$) C (green, $n=22$) D (purple, $n=24$)
average 1.0 1.0 1.0 1.0
complete 0.0 1.0 0.732 1.0
single 0.021 0.0 0.056 1.0
ward 0.0 1.0 0.690 1.0
@@
As mentioned above, cluster D (purple) is visually homogeneous within all four linkage schemes: this is confirmed as its purity scores remain equal to 1. The cells in this cluster have compact shapes and a prominent nucleus, as expected from cells that have been plated on glass. The same types of cells were distinguished as a sub-population FC by Haaster et al. [29]. In contrast, cluster A (red) is much less consistent within the different linkage schemes, with purity scores close to 0 (with the obvious exception of the average linkage). Visually, cells belonging to cluster A are more heterogeneous, with a star-shaped or a triangular shape (first row of Figure 6). This group of cells maps with the sub-population RS identified by Haaster et al.. Cells belonging to cluster B are significantly more elongated. Their purity score is high with the exception of the single linkage scheme, but this could just be anecdotal as there are only 7 cells in this cluster. They may correspond to spindle-shaped elongated cells, fibroblastic in shape, identified as SS cells by Haaster et al.. The cells in cluster C are mostly compact, similar to those in cluster D, but usually larger. The purity scores of cluster C are close to 1, indicating that they form a group with homogeneous shapes. They were probably identified as belonging to the FC subpopulation by Haaster et al..
4.2 $Y1$: heterogeneity of HeLa cell shapes
The set $Y1$ consists of $100$ cells. These cells have already been selected based on manual inspection; see Section 3.4. We followed the procedure described in Section 4.1 for the hMSCs to compute all pairwise distances between the cell contours using the methodology of Section 3.6. The corresponding distance matrix is visualised as a heat map in Figure 7, In contrast to the hMSCs for which we identified a clear outlier (see Figure 3), we do not identify HeLa cells that we would unambiguously classify as outliers.

### v3

@@ Abstract
Quantifying cell morphology is central to understanding cellular regulation, fate, and heterogeneity, yet conventional image-based analyses often struggle with diverse or irregular shapes. We present a computational framework that uses topological data analysis to characterise and compare single-cell morphologies from fluorescence microscopy. Each cell is represented by its contour together with the position of its nucleus, from which we construct a filtration based on a radial distance function and derive a persistence diagram encoding the shape’s topological evolution. The similarity between two cells is quantified using the 2-Wasserstein distance between their diagrams, yielding a shape distance we call the PH distance. We apply this method to two representative experimental systems—primary human mesenchymal stem cells (hMSCs) and HeLa cells—and show that PH distances enable the detection of outliers in those systems, the identification of sub-populations, and the quantification of shape heterogeneity. We benchmark PH against three established contour-based distances (aspect ratio, Fourier descriptors, and elastic shape analysis) and show that PH offers better separation between cell types and greater robustness when clustering heterogeneous populations. Together, these results demonstrate that persistent-homology-based signatures provide a principled and sensitive approach for analysing cell morphology in settings where traditional geometric or image-based descriptors are insufficient.
@@
Cell shape carries important information about how cells grow, respond to their environment, and differ from one another. However, cell populations often contain substantial natural variability, which makes it difficult for traditional image-analysis methods to reliably compare their shapes. In this study, we introduce a new way to measure and compare cell morphology using a mathematical approach known as topological data analysis. Our method describes each cell by the outline of its membrane together with the position of its nucleus, and converts this information into a compact shape signature that can be compared across thousands of cells.
We apply this approach to two widely studied systems—human mesenchymal stem cells and HeLa cells—and show that it can detect unusual cell shapes, reveal sub-populations within heterogeneous samples, and quantify how variable a cell population truly is. We also compare our method with several commonly used shape-comparison techniques and find that it provides clearer separation between distinct cell types and more reliable clustering. Overall, this framework offers a robust and interpretable way to analyse cell morphology and may help researchers better understand the diversity and behaviour of cells in biological experiments.
@@ 1 Introduction
A second representation of 2D shapes, which we refer to as shape as planar contour, is based on the curve describing the outer boundary of the shape (see Figure 1C). This is well suited to applications focused on the geometric configuration of a shape, where factors such as the colour or grey level of the interior are either not relevant or not available. Methods for modelling the similarity between two shapes given as planar contours have been based on defining a distance between two curves in the plane. The proposed distances include the Hausdorff and Frechet distances [8]. Other techniques are based on the Poisson equation [9], integral invariants [10], and an elastic shape distance on the energy required to elastically deform one boundary contour to the other [11, 12, 13, 14].
It is worth providing a little more details on elastic energy as we will use it later for comparison with the new method introduced in this paper. The elastic shape analysis framework, particularly through the Square-Root Velocity Function (SRVF) representation [12], provides a mathematically principled solution by treating shapes as points in an infinite-dimensional Riemannian manifold equipped with an elastic metric. This framework achieves invariance to translation, rotation, scale, and crucially, reparameterization, while allowing for elastic stretching and bending of curves [15, 11]. The SRVF representation transforms the computationally challenging elastic metric into the standard $\mathbb{L}^{2}$ metric, enabling efficient geodesic computation via path-straightening methods and facilitating statistical analyses including mean shape computation, principal component analysis, and shape classification [16, 17]. Recent extensions have broadened the applicability of this framework to medical imaging applications including tumor morphology analysis [18] and fiber tract comparison [15], demonstrating its versatility across diverse scientific domains.
@@
The paper is organised as follows. The next section (Section 2) introduces the concept of persistence homology applied to analysing the morphology of a cell, the construction of the persistence diagram of a cell contour, and the computation of the Wasserstein distance between two persistence diagrams. The Materials and Methods section (section 3) gives information on the experimental data and implementations of the methods mentioned above. It includes a description of three standard methods for computing the distance between two curves (aspect ratio distance, Fourier descriptor distance, and elastic shape distance), as those methods will be compared to our persistence homology distance. The Results section (section 4) discusses the applications of our new method, PH, on two systems, hMSCs that are expected to include sub-populations and, in contrast, HeLa cells whose morphologies are expected to vary continuously. An extensive analysis that compares PH with existing cell contour distances is included. We conclude with a discussion of how to integrate persistent homology with other more traditional techniques for comparing cell shapes.
@@ 2.3 Illustration on a simple cell contour
Readers familiar with persistent homology and persistence diagrams will notice that this is a non-standard modification. Due to the nature of the contours, performing this essential pairing allows us to more efficiently represent and compare the topological descriptors. Further, readers familiar with extended persistent homology will notice that this pairing is baby extended persistence on the boundary curve.
@@ 3.2 HeLa cells
3.3 Fluorescence Staining of Cells
@@ 3.5 Image Processing and Contour Generation
We used the FilamentSensor2.0 tool [28] to perform image processing and extract the contour of each cell from the actin image. Here, we used the feature “Include Area-Outline” to export the contour from the binarised image of the cells. The cell centre (i.e.e the focal point used as a center for the radial function) was defined as the centre of mass of the nucleus, obtained from the aligned microscopy image after segmenting the nucleus in Fiji [29] using Otsu’s thresholding method.
3.6 Converting Pixel Measurements to Physical Units
In the analysis of microscopy images, measurements are first recorded in pixels and need to be translated into physical units, usually microns, to be interpreted accurately. This translation is dependent on the spatial calibration of the imaging system, which is indicated by the pixel size—the actual length each pixel represents. The conversion formula is d = n × s, where d is the size in microns, n denotes the pixel measurement and s is the scale factor in micron per pixel. The scale factor is determined by the total magnification of the objective lens, intermediate optics, and the physical size of the camera sensor pixels. The scale factors are $0.3155$ and $0.1639$ for the $X1$ (i.e., the hMSCs) measurements, and for $Y1$ (i.e., the HeLa cells) measurements, respectively.
@@ 3.8 Clustering cells based on their contour
3.9 Other methods for cell contour comparisons
We compare our method for computing the distance between two cell contours (which we will refer to as PD, for persistence diagram), with traditional distances such as an aspect ratio distance and a distance based on Fourier descriptors, as well as a more recent method based on elastic shape. We briefly describe those three methods below.
3.9.1 A rotation- and translation-invariant aspect ratio distance.
Given a planar contour $C=\{(x_{i},y_{i})\}_{i=1}^{N}$, we first remove translation by centering the points,
$\tilde{C}=C-\frac{1}{N}\sum_{i=1}^{N}(x_{i},y_{i}).$ (3)
Rotation invariance is obtained by applying principal component analysis (PCA) to the covariance matrix of $\tilde{C}$. Let $\lambda_{1}\geq\lambda_{2}$ be the eigenvalues of this covariance matrix. These eigenvalues represent the variances of the contour along the major and minor principal axes, respectively. The (rotation- and translation-invariant) aspect ratio $A$ of the contour is then defined as
$A=\sqrt{\frac{\lambda_{1}}{\lambda_{2}}}\geq 1.$ (4)
An aspect ratio of $A=1$ corresponds to an isotropic shape, while larger values indicate increasing elongation of the contour.
Aspect-ratio distance.
To compare the elongation of two contours with aspect ratios $A_{1}$ and $A_{2}$, we define a distance
$d(A_{1},A_{2})=\left|\log A_{1}-\log A_{2}\right|.$ (5)
The logarithmic transform is preferable to a linear difference $|A_{1}-A_{2}|$ because the aspect ratio is inherently multiplicative: a change from $A=2$ to $A=4$ represents the same geometric deformation as a change from $A=4$ to $A=8$. Using the logarithm therefore produces a scale-symmetric measure of elongation, in which multiplicative changes in aspect ratio correspond to additive differences in the metric. The resulting distance $d(A_{1},A_{2})$ is non-negative, symmetric, and equals zero if and only if the two contours have identical elongation.
3.9.2 A rotation- and translation-invariant distance based on Fourier descriptors.
Fourier descriptors (FDs) provide a compact frequency-domain representation of 2D closed curves. By treating a planar curve as a complex-valued periodic function, FDs capture the shape’s geometric properties in a form amenable to classification and comparison.
Complex Representation of Curves
A closed curve in the plane can be represented parametrically as a complex-valued function:
$z(t)=x(t)+iy(t),\quad t\in[0,1]$ (6)
where $x(t)$ and $y(t)$ are the Cartesian coordinates along the curve, and the parameter $t$ traces the contour from start to finish with $z(0)=z(1)$ (closed curve condition).
For a discrete contour with $N$ sampled points, we have:
$z_{n}=x_{n}+iy_{n},\quad n=0,1,\ldots,N-1$ (7)
Fourier Decomposition
The Fourier descriptors are obtained via the discrete Fourier transform (DFT):
$F_{k}=\frac{1}{N}\sum_{n=0}^{N-1}z_{n}\,e^{-2\pi ikn/N},\quad k=-M,\ldots,M$ (8)
where $k$ is the harmonic index and $M$ is the maximum harmonic considered (typically $M\ll N$).
Note that the original curve can be reconstructed using the inverse DFT:
$z(t)=\sum_{k=-M}^{M}F_{k}\,e^{2\pi ikt}$ (9)
where $t\in[0,1]$ is the normalized curve parameter.
Invariance Properties
To enable meaningful shape comparison, FDs must be invariant to geometric transformations that preserve shape identity:
  • a)
    Translation Invariance
    Translation of the curve by a vector $(a,b)$ adds a constant to $z(t)$:
    $z^{\prime}(t)=z(t)+(a+ib)$ (10)
    This affects only the DC component $F_{0}$. Translation invariance is achieved by:
    1. 1.
      Computing the centroid: $z_{c}=\displaystyle\frac{1}{N}\sum_{n=0}^{N-1}z_{n}$
    2. 2.
      Centering the curve: $z_{n}\leftarrow z_{n}-z_{c}$
    After centering, $F_{0}=0$, and all other coefficients $F_{k}$ ($k\neq 0$) are translation-invariant.
  • b)
    Starting Point Invariance The choice of starting point $t=0$ on the curve is arbitrary. If the starting point shifts by $\tau$, the curve becomes:
    $z^{\prime}(t)=z(t+\tau)$ (11)
    In the frequency domain, this introduces a phase shift:
    $F^{\prime}_{k}=F_{k}\,e^{2\pi ik\tau}$ (12)
    Starting point invariance is obtained by normalizing the phase of all coefficients relative to $F_{1}$:
    $\tilde{F}_{k}=F_{k}\,e^{-ik\theta_{1}},\quad\text{where }\theta_{1}=\arg(F_{1})$ (13)
    This ensures that $\tilde{F}_{1}$ is real and positive, removing the dependence on the starting point.
  • c)
    Rotation Invariance
    Rotation of the curve by angle $\alpha$ multiplies the complex representation by $e^{i\alpha}$:
    $z^{\prime}(t)=e^{i\alpha}z(t)$ (14)
    This affects all Fourier coefficients:
    $F^{\prime}_{k}=e^{i\alpha}F_{k}$ (15)
    Since rotation only changes the phase, rotation invariance is achieved by using the magnitudes:
    $E_{k}=|F_{k}|,\quad k=1,2,\ldots,M$ (16)
The magnitude descriptors $\mathbf{E}=\{E_{1},E_{2},\ldots,E_{M}\}$ are invariant to rotation, translation, and starting point.
An FD distance
To compare two contours $C_{1}$ and $C_{2}$ with magnitude Fourier descriptors $\mathbf{E}^{1}$ and $\mathbf{E}^{2}$, we define a distance
$d(C_{1},C_{2})=\sqrt{\sum_{i=1}^{M}(E_{i}^{1}-E_{i}^{2})^{2}}.$ (17)
The FD distance is parameterized by $M$, the number of harmonics considered.
3.9.3 The elastic distance
The elastic distance, based on the Square-Root Velocity (SRV) framework, provides a mathematically rigorous way to compare shapes while allowing for elastic deformations. Unlike rigid alignment methods, elastic distance accounts for both differences in shape and differences in parameterization, making it particularly suitable for comparing curves that may be non-rigidly deformed versions of each other.
Square-Root Velocity Representation
A planar curve $\mathbf{c}(t)=(x(t),y(t))$, $t\in[0,1]$, is represented by its Square-Root Velocity (SRV) function:
$\mathbf{q}(t)=\frac{\dot{\mathbf{c}}(t)}{\sqrt{\|\dot{\mathbf{c}}(t)\|}}$ (18)
where $\dot{\mathbf{c}}(t)$ denotes the tangent vector. The SRV representation has the key property that the $\mathbb{L}^{2}$ metric on $\mathbf{q}$ corresponds to the elastic metric on the curve $\mathbf{c}$.
The original curve can be recovered from its SRV via integration:
$\mathbf{c}(t)=\mathbf{c}(0)+\int_{0}^{t}\mathbf{q}(s)\|\mathbf{q}(s)\|\,ds$ (19)
Shape Space and Reparameterization
Two curves may represent the same shape but be parameterized differently. A reparameterization is a diffeomorphism $\gamma:[0,1]\to[0,1]$ (smooth, invertible, monotone increasing with $\gamma(0)=0,\gamma(1)=1$). Under reparameterization, the SRV transforms as:
$(\mathbf{q}\circ\gamma)(t)=\mathbf{q}(\gamma(t))\sqrt{\dot{\gamma}(t)}$ (20)
The space of shapes is defined as the quotient space of curves modulo reparameterizations and rigid motions (translation and rotation).
Elastic distance
The elastic distance between two curves $\mathbf{c}_{1}$ and $\mathbf{c}_{2}$ is defined as:
$d_{\text{elastic}}(\mathbf{c}_{1},\mathbf{c}_{2})=\min_{\gamma\in\Gamma,\mathbf{R}\in SO(2)}\|\mathbf{q}_{1}-\mathbf{R}(\mathbf{q}_{2}\circ\gamma)\sqrt{\dot{\gamma}}\|_{\mathbb{L}^{2}}$ (21)
where:
  • $\Gamma$ is the set of all reparameterizations (diffeomorphisms)
  • $\mathbf{R}\in SO(2)$ is a rotation matrix
  • $\|\cdot\|_{\mathbb{L}^{2}}$ is the $\mathbb{L}^{2}$ norm: $\|\mathbf{q}\|_{\mathbb{L}^{2}}^{2}=\int_{0}^{1}\|\mathbf{q}(t)\|^{2}dt$
This minimization finds the optimal alignment (rotation) and reparameterization that brings $\mathbf{c}_{2}$ closest to $\mathbf{c}_{1}$.
Invariance Properties
The elastic distance possesses the following invariances:
  • Translation invariance: Achieved by centering curves at their centroids before computing SRV
  • Rotation invariance: Built into the optimization via $\min_{\mathbf{R}\in SO(2)}$
  • Reparameterization invariance: Built into the optimization via $\min_{\gamma\in\Gamma}$, making it robust to different tracing speeds or starting points
Full descriptions on how to compute the elastic distance is available in Refs [11, 32, 12, 14].
3.9.4 Implementations
We have implemented the computation of the aspect ratio distance and the Fourier descriptor distance as Matlab scripts. The implementation of the elastic energy is a much larger project. Instead of implementing our own version, we relied on the Matlab package fdasrv_MATLAB [33], available at https://github.com/jdtuck/fdasrvf_MATLAB.
3.10 Classical Multidimensional Scaling
Classical multidimensional scaling (cMDS) computes an embedding of a set of $n$ objects into a Euclidean space of dimension $p$ from a matrix of pairwise dissimilarities. Given a symmetric matrix $D=(d_{ij})$ of distances, cMDS assumes that the distances are Euclidean and seeks points $x_{1},\dots,x_{n}\in\mathbb{R}^{p}$ whose pairwise Euclidean distances approximate $d_{ij}$.
Computing the embedding
Let $D$ be the matrix of squared distances, and let $J=I_{n}-\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}$ be the centering matrix. cMDS forms the Gram matrix
$B\;=\;-\tfrac{1}{2}\,J\,D\,J.$ (22)
If the dissimilarities arise from an exact Euclidean configuration, then $B=XX^{\top}$, where $X$ is the matrix whose $i$th row is the coordinate vector of object $i$ in the embedding space.
The matrix $B$ is symmetric and is diagonalized as
$B=V\Lambda V^{\top},$ (23)
where $\Lambda=\mathrm{diag}(\lambda_{1},\dots,\lambda_{n})$ contains the eigenvalues in decreasing order, $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$, and $V$ contains the corresponding orthonormal eigenvectors. Coordinates in $\mathbb{R}^{p}$ are obtained by retaining the largest $p$ positive eigenvalues:
$X_{p}\;=\;V_{p}\Lambda_{p}^{1/2},$ (24)
where $\Lambda_{p}$ is the diagonal matrix of the top $p$ positive eigenvalues and $V_{p}$ the associated eigenvectors (also called principal components).
The contribution or explained variance $C_{p}$ of the principal component $V_{p}$ is defined as:
$C_{p}=\frac{\Lambda_{p}}{\sum_{i=1}^{n}\Lambda_{i}}$ (25)
@@ 4 Results and discussion
The result section serves three purposes:
  • a)
    Identification of subpopulations in a set of cells. We report how our method PH for comparing the shapes of cells can be used to identify unusual cells (i.e., outliers) and subpopulations within a cell population.
  • b)
    Assessing the performance of PH for clustering cells. We check how PH can distinguish two different cell types when the true partitioning is known.
  • c)
    Comparing PH with existing methods for comparing cell contours. We compare PH with an aspect ratio distance, a Fourier descriptor distance, and a elastic distance.
For all three tasks, we consider both human mesenchymal stem cells (hMSCs) and HeLa cells. HeLa cells are a long-established, immortalised cancer cell line. They are a commonly used cell line for culturing in in vitro conditions, showing relatively consistent growth patterns and morphologies within a population, despite some heterogeneity. hMSCs are primary cells, on the other hand, derived from bone marrow directly from patients. They are not immortalised or further altered and exhibit an inherent natural heterogeneity, both in morphology and function, due to their preparation and selection method.
We start with analyzing a population of hMSCs, $X1$ (Section 4.1.1) and a population of HeLa cells, $Y1$ (Section 4.1.2), with the experimental setup and analysis pipeline described in Section 3.
4.1 Finding subpopulations in collections of cells
4.1.1 $X1$: sub-populations of hMSCs
@@
Linkage A (red, $n=86$) B (blue, $n=22$) C (green, $n=7$) D (purple, $n=24$)
average 1.000 1.000 1.000 1.000
complete 0.334 0.008 0.925 1.000
single 0.075 0.025 0.008 1.000
ward 1.000 1.000 1.00 1.000
@@
As mentioned above, cluster D (purple) is visually homogeneous within all four linkage schemes: this is confirmed as its purity scores remain equal to 1. The cells in this cluster have compact shapes and a prominent nucleus, as expected from cells that have been plated on glass. The same types of cells were distinguished as a sub-population FC by Haasters et al. [37]. In contrast, cluster A (red) is much less consistent within the different linkage schemes, with purity scores close to 0 (with the obvious exception of the average linkage). Visually, cells belonging to cluster A are more heterogeneous, with a star-shaped or a triangular shape (first row of Figure 6). This group of cells maps with the sub-population RS identified by Haasters et al.. Cells belonging to cluster B are significantly more elongated. Their purity score is high with the exception of the single linkage scheme, but this could just be anecdotal as there are only 7 cells in this cluster. They may correspond to spindle-shaped elongated cells, fibroblastic in shape, identified as SS cells by Haasters et al.. The cells in cluster C are mostly compact, similar to those in cluster D, but usually larger. The purity scores of cluster C are close to 1, indicating that they form a group with homogeneous shapes. They were probably identified as belonging to the FC subpopulation by Haasters et al..
4.1.2 $Y1$: heterogeneity of HeLa cell shapes
The set $Y1$ consists of $100$ cells. These cells have already been selected based on manual inspection; see Section 3.4. We followed the procedure described in Section 4.1.1 for the hMSCs to compute all pairwise distances between the cell contours using the methodology of Section 3.7. The corresponding distance matrix is visualised as a heat map in Figure 7, In contrast to the hMSCs for which we identified a clear outlier (see Figure 3), we do not identify HeLa cells that we would unambiguously classify as outliers.